2x + 3y = 8
-x + y = -4
Graph the linear system and estimate the solution

Answer by 1998Baby(63)

2x+3y=8
-x+y=-4 => mulitply this whole bottem equation by 2
_______
2x+3y=8
-2x+2y=-8
cross out the x values and add down
you should get 5y=0
divide by five and y =0
then use 0 and put it in any equation.
2x+3(0)=8
2x+0=8
x=4
ordered pair=(4,0)